# PathoTune: Adapting Visual Foundation Model to Pathological Specialists

### v1

@@ Abstract
As natural image understanding moves towards the pretrain-finetune era, research in pathology imaging is concurrently evolving. Despite the predominant focus on pretraining pathological foundation models, how to adapt foundation models to downstream tasks is little explored. For downstream adaptation, we propose the existence of two domain gaps, i.e., the Foundation-Task Gap and the Task-Instance Gap. To mitigate these gaps, we introduce PathoTune, a framework designed to efficiently adapt pathological or even visual foundation models to pathology-specific tasks via multi-modal prompt tuning. The proposed framework leverages Task-specific Visual Prompts and Task-specific Textual Prompts to identify task-relevant features, along with Instance-specific Visual Prompts for encoding single pathological image features. Results across multiple datasets at both patch-level and WSI-level demonstrate its superior performance over single-modality prompt tuning approaches. Significantly, PathoTune facilitates the direct adaptation of natural visual foundation models to pathological tasks, drastically outperforming pathological foundation models with simple linear probing. The code will be available upon acceptance.
@@ 2 Related Work
Pathological Foundation Model. With the emergence of foundation models in natural language processing [28, 3], computer vision [11, 10], etc., recent studies have explored the development of pathological foundation models based on self-supervised learning. Studies like Huang et al. [13] and Ciga et al. [7] apply contrast learning to pathological patches. CTransPath [31] enhances the MoCo v3 framework with a pseudo positive selection mechanism to improve similarity handling between patches. Pathoduet [12] builds on MoCo v3 with additional pretraining tasks for cross-scale and cross-stain challenges. Large-scale data utilization includes HIPT’s [4] hierarchical pyramid ViT pretraining on 10,678 WSI slides, UNI [5] employing 100,000 slides with the DINO v2 framework, and Virchow [29] using 1.5 million slides.
@@ 4 Experiments and Results
Effectiveness of PathoTune. The results of different foundation models using mixed combinations of prompts (Table 1) yield key insights: (1) Our method significantly surpasses linear probing in all scenarios, closely rivals full finetuning with just 5.9% of the trainable parameters. Furthermore, employing multi-modal prompts greatly enhances performance compared to single-modal usage. (2) The performance of a useful downstream adaptation method (e.g., our PathoTune) based on the natural visual foundation model far exceeds that of a poor downstream adaptation approach (e.g., linear probing) based on pathological foundation model, suggesting that efficient downstream adaptation is even more important than pretraining a pathological foundation model. (3) Transferring from a pathological foundation model shows slightly better results than a visual foundation model under the same backbone scale, indicating the optimal strategy originates from a pathological foundation model paired with an effective finetuning approach. (4) PathoTune can significantly enhance underperforming foundation models, e.g., ImageNet (ViT-S), elevating their performance to rival that of specialized models, e.g., Pathoduet (ViT-B).

### v2

@@ Abstract
As natural image understanding moves towards the pretrain-finetune era, research in pathology imaging is concurrently evolving. Despite the predominant focus on pretraining pathological foundation models, how to adapt foundation models to downstream tasks is little explored. For downstream adaptation, we propose the existence of two domain gaps, i.e., the Foundation-Task Gap and the Task-Instance Gap. To mitigate these gaps, we introduce PathoTune, a framework designed to efficiently adapt pathological or even visual foundation models to pathology-specific tasks via multi-modal prompt tuning. The proposed framework leverages Task-specific Visual Prompts and Task-specific Textual Prompts to identify task-relevant features, along with Instance-specific Visual Prompts for encoding single pathological image features. Results across multiple datasets at both patch-level and WSI-level demonstrate its superior performance over single-modality prompt tuning approaches. Significantly, PathoTune facilitates the direct adaptation of natural visual foundation models to pathological tasks, drastically outperforming pathological foundation models with simple linear probing. The code is available at https://github.com/openmedlab/PathoDuet.
@@ 2 Related Work
Pathological Foundation Model. With the emergence of foundation models in natural language processing [28, 3], computer vision [11, 10], etc., recent studies have explored the development of pathological foundation models based on self-supervised learning. Studies like Huang et al. [13] and Ciga et al. [7] apply contrastive learning to pathological patches. CTransPath [31] enhances the MoCo v3 framework with a pseudo positive selection mechanism to improve similarity handling between patches. Pathoduet [12] builds on MoCo v3 with additional pretraining tasks for cross-scale and cross-stain challenges. Large-scale data utilization includes HIPT’s [4] hierarchical pyramid ViT pretraining on 10,678 WSI slides, UNI [5] employing 100,000 slides with the DINO v2 framework, and Virchow [29] using 1.5 million slides.
@@ 4 Experiments and Results
Effectiveness of PathoTune. The results of different foundation models using mixed combinations of prompts (Table 1) yield key insights: (1) Our method significantly surpasses linear probing in all scenarios, closely rivals full finetuning with just 5.9% of the trainable parameters (More details in supplementary material). Furthermore, employing multi-modal prompts greatly enhances performance compared to single-modal usage. (2) The performance of a useful downstream adaptation method (e.g., our PathoTune) based on the natural visual foundation model far exceeds that of a poor downstream adaptation approach (e.g., linear probing) based on pathological foundation model, suggesting that efficient downstream adaptation is even more important than pretraining a pathological foundation model. (3) Transferring from a pathological foundation model shows slightly better results than a visual foundation model under the same backbone scale, indicating the optimal strategy originates from a pathological foundation model paired with an effective finetuning approach. (4) PathoTune can significantly enhance underperforming foundation models, e.g., ImageNet (ViT-S), elevating their performance to rival that of specialized models, e.g., Pathoduet (ViT-B).
@@ 5 Conclusion
{credits}
5.0.1 Acknowledgements
This study was supported by Shanghai Artificial Intelligence Laboratory.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.